# LLM-Resistant Math Word Problem Generation via Adversarial Attacks

### v1

@@ 4.3 Number of Adversarial Examples
While on average each problem generates over a thousand modifications, it is worth noting that the actual number of possible generations varies significantly based on the values and the number of variable nodes in a problem. For example, a problem with only one digit value and one variable node has a potential maximum of 9 possible modifications under the $M3$ generation method. Given this reason and computational cost considerations, in this work we randomly select 100 problems from MultiArith and GSM8K separately and generated 100 adversarial examples for each problem. Despite the seemingly small number of problems, the creation of 30,000 adversarial examples per dataset given three different generation methods (totaling 60,000 examples across both datasets) is significant. We present the GSM8K results in Figure 3, observing a clear pattern of accuracy decline across all models as the number of examples increases. Interestingly, for several models, just 10 adversarial examples are enough to degrade the performance.
4.4 Main Results
@@
We evaluate each adversarial example for each problem against all models given different generation methods. A problem is considered incorrect if there exists at least one adversarial example under that problem that results in an incorrect answer. We refer such problem as successful attack. Here we do not run the full generated adversarial examples against GPT-4 due to cost consideration. Instead, we propose a cost-effective way to query expensive models in §4.6. We report the model performance with all three generation methods in Table 1. Ranging from the most lenient method, $M1$, where the numbers could be fairly wild, to the most stringent condition, $M3$, we observe a consistent accuracy drop across all models, even for the strong models (e.g., math-tuned, larger size, or API-based). For the most original-problem-like generation, $M3$, we observe that the weaker models (e.g., smaller size or general-purpose) fail to generate any correct answers in every case. As a comparison to rephrasing attack, we present the average ASR of $M3$ in Table 2. Even as the most restrictive generation method, $M3$ significantly outperforms RobustMath in every model.
Model ASR (RobustMath) ASR (ours)
Mistral 7B 0.0 100.0
MetaMath 7B 13.0 48.5
Llama-2 13B 0.0 100.0
WizardMath 13B 1.0 82.5
Vicuna 13B 0.0 97.5
CodeLlama 34B 67.1 100.0
MetaMath 70B 11.1 31.0
GPT-3.5-Turbo 16.9 33.5
Table 2: Comparison Result: given that RobustMath comprises MultiArith and GSM8K, we calculat the average ASR of $M3$ from our results for MultiArith and GSM8K. Our method significantly outperforms RobustMath in every model.
4.5 Universal Attacks
We are interested in whether there exist problems that all LLMs fail, also known as universal attacks. We count the number of successful attacks from each model and calculated the percentage of common successful attacks among them. We present the results in Table 3 and observe a clear pattern of decreasing universal attacks with an increased number of models. This suggests that diversity among models can lead to increased resilience against such attacks. We also observe that the impact of universal attacks varies significantly across different generation methods, with $M3$ having the lowest number of universal attacks.
@@ 4.5 Universal Attacks
GPT-3.5-Turbo 8 9.0 46.0 83.0
Table 3: Universal Attack: the number of universal attacks decreases as the number of models increases, implying that diversity among models enhances resilience against such attacks.
4.6 Efficient Attacks
@@
Figure 3 shows that scaling up the number of adversarial examples results in a steady performance drop for all models. However, this scaling might not be viable for API-based models such as GPT-4 or GPT-3.5 due to request limits and high costs. To circumvent this, we propose an efficient attack method for API-based models by using adversarial examples in a targeted manner. The idea is simple: we attack model $A$ (target model, GPT-4 in our case), with successful attacks from a cheaper model $B$ (e.g., open-source, lower API cost). We first select 50 problems along with their 100 adversarial examples running against GPT-4. Next, we select 50 successful attacks from a cheaper model and run them against GPT-4. Note that for some models, there are no correct responses, even with $M3$. Therefore, we only compare $M3$ with the models that has correct answers. We compare the results in Table 4 and observe a significant request reduction while achieving similar ASR.
Method Req. RR. (%) ASR (%)
Original 5000 0 10
MetaMath 7B 1389 72 8
WizardMath 13B 1745 65 6
MetaMath 70B 672 87 8
GPT-3.5-turbo 456 91 10
@@
Table 4: Efficient Attack: a targeted approach to attack high-cost models. We leverage successful attacks from cheaper models to attack GPT-4, achieving the same ASR with about 90% reduction in requests.
4.7 Human Evaluation
We conduct a human evaluation on the generated adversarial examples to ensure they are indeed valid. We randomly select one adversarial example from 20 problems from all three generation methods (totaling 60 examples) that GPT-3.5 failed on for GSM8K and presented them to three human evaluators to assess (i) correctness: whether the answer to the modified problem is correct; (ii) coherence: whether the problem is contextually coherent; and (iii) similarity: whether the newly generated problem’s difficulty level is similar to the original one. The annotators were asked to make a binary decision to determine whether the criteria described above were met. We present the average of their scores in Figure 4. Overall, we found that all human evaluators agreed with the correctness. $M3$ scores highly across all metrics, indicating that it indeed preserves the original problem’s coherence and difficulty level. On the other hand, the scores for coherence and similarity varies between generation methods, suggesting that while they all successfully perturbed the model, some of the examples may not be useful in an educational context.
@@ 4.7 Human Evaluation
Figure 4: Human Evaluation: we report the average score among three annotators on the successful attacks from GPT-3.5-turbo. $M3$ shows the highest score across all metrics, indicating our best generation method correctly generate contextual coherent problems that preserve the original level of difficulty.
@@ 5 Analysis and Discussion
In this section we conduct several analyses to better understand our generation methods and LLMs mathematical capabilities.
5.1 Transferability
@@
We investigate the transferability of adversarial examples among the models. We attack model $A$ with the successful attacks from model $B$ and calculate the number of common successful attacks between these two models. We present the GSM8K $M3$ result in Figure 5. Interestingly, we observe that Vicuna 13B, Llama2 13B, CodeLlama 34B, and Mistral 7B exhibit a high percentage of transferability, reaching 100%, indicating significant vulnerability and a strong correlation among these models. On the other hand, comparisons with MetaMath 70B and GPT-3.5 tend to have a lower transferability rate. While it might suggest a form of resistance to adversarial examples that affect other models, it is not a measure of robustness as it could be that those models fail on entirely different examples. This finding suggests that certain LLMs might struggle with particular numbers or patterns. To further understand these phenomena, we conduct further analysis in §5.2.
Figure 5: Transferability: we present the adversarial example transferability amount all models by comparing each model against all other models. Compare to the math-tuned and production models, the weaker models such as Mistral 7B and LLama2 13B exhibit significant vulnerability and a strong correlation amount them.
5.2 Question Feature Analysis
As a step towards understanding problems that are hard for LLMs, we conduct regression analysis to understand the effects of various features of the problems. More specifically, we construct features from the 20,000 MWPs using the generation method $M3$ (200 problems, each with 100 modifications). The input features include operation counts, value ranges, and node counts in the AST, resulting in a total of 51 input features. The features are plotted against the correctness of the model, with correct being $1$ and wrong being 00.
We present the coefficients of the eight models with some selected features in Table 7 in the Appendix. The coefficients are only compared within each model, but not across models. A more suited metric to compare between models is accuracy, such as Table 1 and Figure 3. We present a set of graphs showing models’ accuracy versus some prominent features in Figure 7. Overall, we found that the value of the answer has a stronger correlation to the correctness. Somewhat surprisingly, most models have positive correlations between the number of constants and variables in a question and the correctness. The reason might be that the problems with more numbers have more straightforward operations (additions and subtractions). Additionally, we found that most high-performing models exhibit a negative correlation with the number of divisions in the problem, except for MetaMath 70B, and operation counts negatively correlate with correctness for all models.
5.3 Code Analysis
@@
While the majority of generated Python code adheres to correct problem-solving logic and produces answers correctly, to ensure that the generated code is correctly being converted into ASTs, we manually validate the generated ASTs and identify the following errors: (1) Incorrect answers or code; (2) Use of loops, conditional, and comparison statements, and functions; (3) Use of expressions such as x = y // z + (y % z > 0); (4) Number misalignment, where the same number appears multiple times in the code or the problem. For GSM8K, 7 out of 200 problems were filtered out. For MultiArith, 15 out of 200 problems were filtered out. We discuss the methods of identifying and handling such errors in Appendix F.
6 Conclusion
@@
In conclusion, we introduce a novel approach to generate MWPs that challenge the LLM’s math-solving abilities, thereby addressing the urgent need for methods that ensure fair educational evaluations. By utilizing AST to create adversarial examples, our generation method significantly degrades both open- and closed-source LLMs without compromising the problems’ original difficulty and coherence. We investigate the shared mathematical vulnerabilities among LLMs and propose a cost-effective method to attack high-cost API-based models, achieving up to a 90% decrease in request rates without compromising performance. Our automatic analyses shed light on the specific mathematical aspects that most influence LLM performance. This work not only paves the way for further research into developing more robust and ethically sound educational tools but also offers valuable insights into LLMs’ limitations, contributing to the ongoing discourse on the ethical use of LLMs in education.
7 Limitations
@@
Perceived Difficulty Level: In this work, the correlation between the complexity of problems generated to challenge LLMs and the actual difficulty perceived by human students has not been empirically validated. It is plausible that modifications considered neutral or simple for LLMs could be perceived as more challenging by human students. We plan to further investigate this in future work.
Imperfect Logical Coherence: While $M3$ produces similar numbers to the original problems, logical coherence is not perfect in few cases. For example, the phrases like “25 hour in a day” and “37 days in a month”. While this issue can be easily addressed by employing heuristic rules to associate certain numbers (e.g., 7, 24, 365) with specific entities (e.g., “days”, “hours”, “year”), we try to keep everything as generalized as possible without overfitting to specific cases.
Cannot Alter Words: In this work, we only consider actual numeric values in a question; however, there are problems where words also represent numeric values. For example, a question from GSM8K is “A frog lays her eggs over a series of 4 days. The first day she lays 50 eggs. The second day, she doubles her production of eggs. The third day she lays 20 more than the second day, and the last day she doubles the first three days total. How many eggs did the frog lay over the span of the 4 days?” contains words like “doubles” and “three” that also represent numeric values. In the future, we plan to add a preprocessing step for the problems to extract as many numeric values from a question as possible, resulting in more diverse generations of problems.
@@ 8 Ethics Statement
9 Acknowledgements
We are very grateful to Sandy Yeung and Jack Goffinet for the helpful initial project discussion. Thank you to the anonymous reviewers for their feedback. The research reported here was supported by the Learning Engineering Virtual Institute funded by leading education philanthropists and organizations through Grant G-23-2137070, to the University of Florida and its partner institutions. The opinions expressed are those of the authors and do not represent the views of the University of Florida, the partner institutions, or those of the philanthropists and organizations.
Appendix A Prompt Template
@@ A.1 Code Generation Prompt
[…]
@@ A.2 Zero-shot CoT
[…]
@@ Appendix C AST Nodes
We associate the number in the original question with the variable node so that a change in the variable node will also change the number in the question. If a number appears multiple times in the code and the problem, then we associate them based on the specific procedures described in Section F.4.
@@ Appendix D Generation Details
                     Let $d$ be the number of digits $val(v_{i})$ has. if $d=1$ then
@@
$val(v^{\prime}_{i})\leftarrow a^{\prime}_{i}\times 10^{b}$ return $v^{\prime}_{i}$
@@
$attempts\leftarrow 0,j\leftarrow 1$ $W^{\prime}=\emptyset$ while $attempts<N$ do
      $V^{\prime}_{j}=\emptyset$ for $v_{i}\in V$ do
           Obtain $v^{\prime}_{i}$ from Algorithm 1 with input $v_{i}$, $M$, and $c$ $V^{\prime}_{j}\leftarrow V^{\prime}_{j}\cup\{v^{\prime}_{i}\}$
           $attempts\leftarrow attempts+1$ $S^{\prime}\leftarrow S$ Apply the new sequence of values from the variable nodes $V^{\prime}_{j}$ to $S^{\prime}$ $accept\leftarrow True$ for each node $s^{\prime}_{k}$ in $S^{\prime}$ do
@@
                          $W^{\prime}\leftarrow W^{\prime}\cup\{V^{\prime}_{j}\}$ $j\leftarrow j+1$
@@
Original
Sarah makes 5 times more money per hour than
Connor does. If Connor earns \$7.2 per hour, how
much does Sarah make in an 8-hour day?
288
M3
Sarah makes 7 times more money per hour than
Connor does. If Connor earns \$8.0 per hour, how
much does Sarah make in an 7-hour day?
392
M2
Sarah makes 20 times more money per hour than
Connor does. If Connor earns \$7.7 per hour, how
much does Sarah make in an 8-hour day?
1232
M1
Sarah makes 76780 times more money per hour than
Connor does. If Connor earns \$7655.5 per hour, how
much does Sarah make in an 12568-hour day?
7387335796720
@@
Table 5: Three examples of questions generated by the three generation methods.

### v2

@@ 4.2 Our Attacks
We set the number of attempts to generate problem variants to be 30,000 and present the average number of generated variants per problem for each $M$ in Appendix G. While on average each problem generates over a thousand variants, it is worth noting that the actual number of possible generations varies based on the original values and the number of variable nodes in a problem. [^7] For example, a problem with only one digit value and one variable node has a potential maximum of 9 possible modifications under $M3$. Given this reason and cost considerations, in this work we equally select 100 random problems from MultiArith and GSM8K and generate 100 variants for each selected problem in each $M$. Despite the seemingly small number of selected problems, the total data points are up to 60,000.8883 generation methods $\times$ 200 selected problems $\times$ 100 variants = 60,000.
Main Attacks
@@
We report the model performance from all $M$s in Table 1.999We do not run the full dataset against GPT-4 due to cost considerations. Instead, we propose a cost-effective approach to query expensive models in Efficient Attack section. Ranging from the most lenient method, $M1$, where the numbers could be fairly wild, to the most stringent one, $M3$, we observe a consistent accuracy drop across all models, even for the strong models (e.g., math-tuned, larger size, or API-based). For the most original-problem-like generation, $M3$, we observe that weak models (e.g., smaller size or general-purpose) fail to generate any correct answers in every case. The left of Table 2 shows each model’s average percentage of incorrect variants. It is worth noting that these percentages are highly correlated with the reported ASR, with GPT-3.5 and MetaMath being the most robust, and CodeLLama and LLama 2 being the least robust models. Furthermore, we compare $M3$ with rephrasing attacks in Table 3. Our method significantly outperforms the baseline in every model, resulting in 60.5 ASR point improvement on average. We also analyze how ASR changes given different numbers of attacks in Appendix H. Interestingly, for several models, just 10 attacks are enough to degrade the performance significantly.
Models
AE. (%)
Mistral 7B 70.5
MetaMath 7B 15.2
Llama 2 13B 91.6
WizardMath 13B 21.0
Vicuna 13B 50.3
CodeLlama 34B 91.8
MetaMath 70B 7.9
GPT-3.5 6.9
@@
GPT-3.5 8 9.0 46.0 83.0
Table 2: Incorrect Variants (Left) The average percentage of incorrect variants per problem for each model. A detailed count distributions can be found in Appendix F. Universal Attack: (Right) All models share some percentage of universal attacks. Increasing the number and diversity among models could enhance their robustness against such attacks.
RobustMath Ours ($M3$)
Model OA AA ASR OA AA ASR $\Delta$ ASR
Mistral 7B 10.3 18.7 0.0 33.0 0.0 100.0 +100.0
MetaMath 7B 91.1 79.3 13.0 97.5 51.0 48.5 +35.5
Llama-2 13B 2.3 8.3 0.0 11.0 0.0 100.0 +100.0
WizardMath 13B 71.0 70.3 1.0 89.0 15.5 82.6 +81.6
Vicuna 13B 46.3 51.7 0.0 68.0 2.0 97.5 +97.5
CodeLlama 34B 31.3 10.3 67.1 8.5 0.0 100.0 +32.9
MetaMath 70B 93.0 82.7 11.1 98.5 68.0 31.1 +20.0
GPT-3.5 91.1 75.7 16.9 94.0 63.0 33.3 +16.4
Average 54.6 49.6 13.6 62.4 24.9 74.1 +60.5
Table 3: Comparison Result We calculate the average of each metric from both datasets and compare the result with rephrasing attack, RobustMath. Our method significantly outperforms the baseline in every model, with an average improvement of 60.5 ASR points.
Universal Attacks
Following the previous adversarial attack literature, we investigate whether there are adversarial examples exist in all LLMs, also known as universal attacks (Zou et al., 2023; Moosavi-Dezfooli et al., 2017). We count the number of adversarial examples from each model and calculated the percentage of common ones among them. We report the results on the right of Table 2 and observe a clear pattern of decreasing universal attacks with an increased number of models. This suggests that while all models are vulnerable to some degree of universal attacks, the percentage of such attacks decreases as more and diverse models are considered. We also observe that universal attacks varies significantly across different $M$s, with $M3$ having the lowest number of universal attacks.
Efficient Attacks
@@
Scaling up the number of attacks results in a steady performance drop for all models. However, this scaling might not be viable for API-based models such as GPT-4 due to request limits and high costs. To circumvent this, we propose an efficient attack method for API-based models by using adversarial examples in a targeted manner. The idea is simple: we attack model $A$ (target model, GPT-4 in our case), with adversarial examples from a cheaper model $B$ (e.g., open-source, lower API cost). We select 50 problems along with their 100 variations and run them against GPT-4, comparing the results with attacking GPT-4 using 50 adversarial examples from a cheaper model. Note that for some models, there are no correct responses, even with $M3$. Therefore, we only compare $M3$ with the models that have correct answers. We compare the results in Table 4 and observe a significant request reduction while achieving similar ASR.
Models Request Call Cost (\$) ASR $\Delta$ Request (%)
MetaMath 7B 1,389 10.7 8.0 72.2
WizardMath 13B 1,745 8.8 6.0 65.1
MetaMath 70B 672 4.3 8.0 86.6
GPT-3.5 456 2.8 10.0 90.9
Target: GPT-4 5,000 29.3 10 -
@@
Table 4: Efficient Attack A targeted approach to attack high-cost models. We leverage adversarial examples from cheaper models to attack GPT-4, achieving the same ASR while reducing up to 90% request calls.
@@
Figure 2: Human Evaluation (Left) We report the average score among three annotators on the adversarial examples from GPT-3.5. $M3$ has the highest score across all metrics, indicating our best generation method correctly generates contextually coherent problems that preserve original difficulty. Transferability (Right) We present the adversarial example transferability among all models by comparing each model against all other models. Compared to the math-tuned and production models, the weaker models such as Mistral 7B and LLama2 13B exhibit significant vulnerability and a strong correlation among them.
@@ 5 Analysis and Discussion
In this section, we conduct analyses and discussions to better understand our generation methods and the mathematical capabilities of LLMs.
Human Evaluation
We conduct a human evaluation to ensure the our generation are indeed valid. Three human evaluators are asked to assess (i) correctness: whether the answer to the modified problem is correct; (ii) coherence: whether the problem is contextually coherent; and (iii) similarity: whether the newly generated problem’s difficulty level is similar to the original one. We randomly select one adversarial example from 20 problems that GPT-3.5 failed on for GSM8K. The annotators made binary decisions to determine whether the criteria were met. We present the average of their scores on the left of Figure 2. Overall, all human evaluators agree on the correctness. $M3$ scores highly across all metrics, indicating that it preserves the original problem’s coherence and difficulty level. Coherence and similarity are relatively low for $M2$ and $M3$, suggesting that while they might perturb the model, such generation may not be useful for the educational context.
Transferability
@@
We investigate the transferability of adversarial examples among the models. We attack model $A$ with the adversarial examples from model $B$ and calculate the number of common adversarial examples between these two models. We present $M3$ result in the left of Figure 2. Interestingly, we observe that weaker models exhibit a high percentage of transferability, indicating a strong vulnerability correlation among these models. On the other hand, comparisons with strong models such as MetaMath 70B and GPT-3.5, tend to have a lower transferability rate. While it might suggest a form of resistance to adversarial examples that affect other models, it is not a measure of robustness as it could be that those models fail on entirely different examples. This finding suggests that certain LLMs might struggle with particular numbers or patterns. To further understand these phenomena, we conduct regression analysis as shown below.
Question Feature Analysis
To further understand the limitations of LLMs, we conduct a regression analysis to investigate the effects of various features of MWPs on the correctness of the model’s predictions. We construct a total of 51 input features from 20,000 $M3$ generated problems, including features such as operation counts, value ranges, and node counts in the AST. The features are plotted against the correctness of the model, with correct being 1 and wrong being 0. We present the coefficients of the all LLMs with some selected features in Table 7 in the Appendix. We observe that the most positively and negatively correlated features vary across different models, suggesting that our adversarial examples exploit different weaknesses of each model. For example, models exhibit varying performance with different answer value ranges. Mistral 7B and WizardMath 13B perform relatively well on problems with smaller answer values (e.g., in the range of [2, 8) and [8, 32)), while MetaMath 7B and Vicuna 13B show better performance on problems with answer values in the range of [32, 128). This suggests that different models may have learned to specialize in problems with specific range, possibly due to the varying distributions of the pre-training data. Overall, problems involving division and more operations tend to be more challenging for most LLMs. This confirms the intuition that complex problems requiring more mathematical operations are generally more difficult for LLMs to solve (Imani et al., 2023b). Visualizations are available in Figure 7 in the Appendix.
Python Code Analysis
@@
While the majority of generated Python code adheres to correct problem-solving logic and produces answers correctly, to ensure that the generated code is correctly being converted into ASTs, we manually validate the generated ASTs and identify the following four types of possible errors: (1) Incorrect answers or code; (2) Use of loops, conditional, and comparison statements, and functions; (3) Use of expressions such as x = y // z + (y % z > 0); (4) Number misalignment, where the same number appears multiple times in the code or the problem. Encouragingly, for GSM8K and MultiArith, 96.5% and 92.5% of ASTs, respectively, are found to be valid and free from these errors. We discuss the methods of identifying and handling such errors in Appendix I.
Why LLMs suffer From Such Simple Attack?
While LLMs can be trained on vast amounts of math data, they may not have an inherent grasp of the underlying mathematical concepts and reasoning steps (Saxton et al., 2019; Lample & Charton, 2019). Instead, they likely rely on pattern and statistical associations learned from the training data to make predictions (Bender et al., 2021), which is also know as memorization (Carlini et al., 2022). Modifying the numbers in a problem may disrupt these learned patterns, causing the models to make errors. This is evidenced by the significant performance drops observed in models like WizardMath (89% to 20%) and Vicuna (76% to 4%) when presented with adversarial examples. Furthermore, test data contamination could also play a role - if the training corpus contains the test data, the models may be able to answer the original questions correctly but struggle to generalize to even subtle variations (Balloccu et al., 2024).
6 Conclusion and Future work
@@
We introduce a novel approach to generate MWPs that challenge the LLM’s math-solving abilities, thereby addressing the urgent need for methods that ensure fair educational evaluations. By utilizing AST to create adversarial examples, we preserve problems’ original difficulty and coherence while significantly degrading both open- and closed-source LLMs and surpassing the previous rephrasing attack by an average of 60% ASR. We conduct a human evaluation to validate the adversarial examples and investigate universal attacks and attack transferability, proposing a cost-effective method to attack high-cost API-based models, achieving up to a 90% request reduction while maintaining the same performance. Our automatic regression analyses demonstrate that our adversarial examples exploit different weaknesses of each model, shedding light on the specific mathematical aspects that most influence individual LLM performance. This work not only paves the way for further investigation into the mathematical limitations among LLMs but also contributes to developing more robust and ethically sound educational tools to ensure fair evaluation, contributing to the ongoing discourse on the ethical use of LLMs in education.
@@
In this work, the correlation between the complexity that challenges LLMs and the actual difficulty perceived by humans has not been empirically validated. In future work, we aim to explore this and expand the methods to handle more complex problems.
@@ 7 Ethics Statement
Appendix A Prompt Templates
@@ A.1 Code Generation Prompt
Code Generation Prompt Write a Python script that contains a step-by-step solution for a given math problem and its answer. Start the script by defining variables corresponding to the quantities mentioned in the problem. Print the final answer at the end of the script. Do not use if-else statements, floor divisions, and loops to solve the problem. Given problem: {problem} Answer: {answer} Python script:
@@ A.2 Zero-shot CoT
Evluation Prompt Solve a math problem. The solution ends with “the answer is (a number)” like “the answer is 1”. Question: {problem} Answer: Let’s think step by step.
@@ Appendix C AST Nodes
We associate the number in the original question with the variable node so that a change in the variable node will also change the number in the question. If a number appears multiple times in the code and the problem, then we associate them based on the specific procedures described in Section I.4.
@@ Appendix D Generation Details
Let $a_{i}\times 10^{b}_{i}\approx h=val(v_{i})$, where $a_{i}$ is an integer that is not divisible by 10 and has at most $c$ scientific digits.
@@
                     Let $d$ be the number of digits $val(v_{i})$ has.
                     if $d=1$ then
@@
                                              $val(v^{\prime}_{i})\leftarrow a^{\prime}_{i}\times 10^{b}$
return $v^{\prime}_{i}$
@@
$attempts\leftarrow 0,j\leftarrow 1$
$W^{\prime}=\emptyset$
while $attempts<N$ do
      $V^{\prime}_{j}=\emptyset$
      for $v_{i}\in V$ do
           Obtain $v^{\prime}_{i}$ from Algorithm 1 with input $v_{i}$, $M$, and $c$
           $V^{\prime}_{j}\leftarrow V^{\prime}_{j}\cup\{v^{\prime}_{i}\}$
           $attempts\leftarrow attempts+1$
           $S^{\prime}\leftarrow S$
           Apply the new sequence of values from the variable nodes $V^{\prime}_{j}$ to $S^{\prime}$
           $accept\leftarrow True$
           for each node $s^{\prime}_{k}$ in $S^{\prime}$ do
@@
                          $W^{\prime}\leftarrow W^{\prime}\cup\{V^{\prime}_{j}\}$
                          $j\leftarrow j+1$
@@
Table 5: Examples of questions generated by the three generation methods.